OpenStudy (anonymous):

Find all solutions in the interval [0,2pi) for sin2x + sqr(3)cosx = 0

OpenStudy (anonymous):

\[Sin2x + \sqrt{3} cosx=0\]

OpenStudy (mertsj):

replace sin2x with 2sinxcosx and factor out cosx

OpenStudy (anonymous):

so it would be \[cosx (2sinx + \sqrt{3})\]? @Mertsj

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Now solve each factor separately. \[\cos x=0\]and\[\sin x=\frac{-\sqrt{3}}{2}\]

OpenStudy (anonymous):

thanks :) i finished